# CodeScore: Evaluating Code Generation by Learning Code Execution

### v3

@@ Abstract
1 Introduction
@@
Automatic evaluation of code generation is significant and promising in the fields of natural language processing (NLP) and software engineering. Due to the great potential of code generation in reducing development costs and revolutionizing programming modes, both industry and academia have devoted substantial attention to it (Li et al., 2022; Mukherjee et al., 2021; Yin and Neubig, 2018; Chen et al., 2021; Shen et al., 2022; Dong et al., 2022). Code generation has achieved remarkable developments in the past few years (Fried et al., 2022; Nijkamp et al., 2022; Dong et al., 2023b; Jiang et al., 2023), but CEMs still need to catch up. It is challenging to evaluate the competitiveness of various approaches without proper CEM, which hampers the development of advanced techniques for code generation. A range of code generation subtasks would benefit from valid code evaluation, including code completion (Guo et al., 2022b; Lu et al., 2022), code translation (Rozière et al., 2020; Zhu et al., 2022), code search (Sun et al., 2022; Arakelyan et al., 2022), etc. Therefore, research on code evaluation is necessary and should be put on the agenda.
@@
Figure 1: Results of evaluating the generated code implementing bubble sort using different CEMs. BLEU and CodeBLEU score the truly functional correct code (c) lower than the incorrect code (b).
@@
Some commonly used match-based CEMs treat code as text, such as BLEU (Papineni et al., 2002) and Accuracy, which focus on basic and lexical-level features. They compute scores mainly based on n-gram co-occurrence statistics. CodeBLEU (Ren et al., 2020) additionally takes into account the structure of code, i.e., abstract syntax tree and data flow. However, the preceding CEMs have deficiencies in identifying code relationships, because code is mainly evaluated based on functional correctness rather than exact/fuzzy match to reference code, and match-based CEMs cannot account for the large and complex space of code functionally equivalent to reference code (Inala et al., 2022). For example, in Fig. 1, code (a) and code (b) have a much higher similarity of tokens or structures than code (c). But through execution, we realize that code (a) and code (c) are different renderings of the same function. By contrast, the execution result of code (b) differs dramatically from both other codes, and code (b) even fails to compile. As a result, merely measuring the similarity of token/structure is insufficient for code evaluation.
@@
LLMs pre-trained on code have demonstrated outstanding results in code generation tasks (Chen et al., 2021; Fried et al., 2022; Li et al., 2022; Dong et al., 2023a), which are fundamentally dependent on exceptional code comprehension. Excellent code comprehension is a crucial element for facilitating code evaluation. We hypothesize that LLMs pre-trained on code possess the ability to evaluate code. However, due to the training strategy of predicting the next token according to context, they lack awareness of evaluating code for functional correctness. Our objective is to instruct LLMs to evaluate code effectively in terms of functional correctness.
@@
Another issue that requires resolution is that the existing match-based CEMs are exclusively confined to the Ref-only input format. This restriction presents three inherent disadvantages. First, for any code generation task, the correct solutions are not finite, but rather, they are inexhaustible. In this context, the provided reference code merely represents one correct solution among a vast multitude. Therefore, it is overly narrow to compare the generated code solely with one correct solution. Second, they neglect the natural language (NL) description, which is a rich repository of information and a real requirement source. Third, these metrics are unusable in the absence of a reference code. This situation is quite commonplace in real-world evaluations where a correct solution is not always readily available. As a result, expanding the input format of CEM is necessary.
@@
In this paper, we propose an effective LLM-based CEM, called CodeScore, which measures the functional correctness of generated codes on three input formats (Ref-only, NL-only, and Ref&NL). To obtain CodeScore, we present a code evaluation learning framework, UniCE, for tuning LLMs to estimate execution similarities with unified input. Specifically, we finetune LLMs to learn PassRatio and Executability of generated code, where Executability is devised to distinguish between compilation errors and output errors for code with PassRatio equal to 0. Generally, codes exhibiting higher functional correctness will pass more test cases, thereby achieving a higher PassRatio  [^1] Note that, although PassRatio varies across different test cases, it tends to yield a higher PassRatio for high-quality code, since we generate a large number of test cases. This phenomenon is somewhat akin to the process of human feedback. Despite the inherent variability in scores assigned by different human evaluators, the overarching trend remains consistent.. Consequently, for unexecutable codes, the model tends to assign scores approaching zero. In contrast, for codes demonstrating superior functional correctness, the model is likely to assign higher scores. CodeScore has the following advantages: 1) CodeScore has excellent evaluation performance, which achieves the state-of-the-art performance correlation with functional correctness on multiple code evaluation datasets. 2) CodeScore provides three application scenarios (Ref-only, NL-only, and Ref&NL) for code evaluation with unified input, while traditional CEMs only consider Ref-only. Our major contributions can be summarized as follows:
@@
  • We propose an efficient and effective LLM-based CEM, CodeScore, that accommodates the functional correctness of generated codes from execution viewpoint. [^2] https://huggingface.co/dz1/CodeScore
@@
  • CodeScore substantially outperforms match-based CEMs and LLM-based EMs, and achieves the state-of-the-art performance on multiple code evaluation datasets.
Figure 2: Examples of three input formats for code evaluation.
2 Methodology
@@ 2.1 CodeScore
$\operatorname{PassRatio}=\frac{1}{|C_{p}|}\sum_{c\in C_{p}}\mathbb{I}\left\{\operatorname{Eval}\left(\mathbf{g}_{p},\mathcal{I}_{p,c}\right)=\mathcal{O}_{p,c}\right\}.$ (1)
@@
$\operatorname{Executability}=\left\{\begin{aligned} &1,\ if\ code\ is\ executable,\\ &0,\ otherwise.\end{aligned}\right.$ (2)
Figure 3: Diagram of UniCE, where the left side of the figure shows its model architecture, and the right side of the figure shows the example (case in Fig. 1) of input and output.
@@
$\operatorname{CodeScore},\operatorname{Exec}=\operatorname{UniCE}(\mathbf{x}),$ (3)
@@
We encourage UniCE to learn code execution (i.e., PassRatio and Executability) by minimizing loss function $\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{E}$:
$\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{E},$ (4)
$\mathcal{L}_{C}=\left(\operatorname{CodeScore}-\operatorname{PassRatio}\right)^{2},$ (5)
$\mathcal{L}_{E}=-\log\mathbf{p}(\operatorname{Exec}|\operatorname{Executability}),$ (6)
where
$\mathbf{p}(\operatorname{Exec}|\operatorname{Executability})=\left\{\begin{aligned} &\mathbf{p}(\operatorname{Exec}),\ if\ \operatorname{Executability}=1,\\ &1-\mathbf{p}(\operatorname{Exec}),\ otherwise.\end{aligned}\right.$ (7)
2.2 UniCE
@@
2.2.1 Pooling Layer
. For LLMs, the pooling layer plays a critical role in enhancing the model’s ability to capture and utilize information more effectively. The work (Tenney et al., 2019; Zhang et al., 2020; Rei et al., 2020) shows that exploiting information from different layers of LLM generally results in superior performance than only the last layer. Therefore, following the work (Peters et al., 2018), we pool information from different layers by using a layer-wise attention mechanism, and the final embedding of a token $t$ can be computed as:
$e_{t}=\gamma\sum^{l}_{k=1}e^{k}_{t}h^{k},$ (8)
@@ 2.2.1 Pooling Layer
2.2.2 Unified Embedding
. We require an efficient and comprehensive representation to encapsulate the unified input sequence $x$. Generally, there are two standard methods to extract the representation of $x$, i.e., averaging all token embeddings and using the first token embedding. While the first method is straightforward and includes information from all tokens, it may dilute the significance of more critical tokens and introduce extraneous noise. Moreover, the work (Ranasinghe et al., 2020; Wan et al., 2022) also proves the superiority of using the first token embedding compared to averaging all token embeddings in various applications. Thus, we employ the final embedding of first token $e_{first}$ as the representation of unified input sequence $x$.
2.2.3 Unified training
. In UniCE, $e_{first}$ is fed to a feedforward neural network to output a score and/or a category. To unify three evaluation input formats into UniCE, we apply multi-task learning for training. Specifically, for each step, we assign three sub-steps for three input formats, yielding $\mathcal{L}^{Ref}$, $\mathcal{L}^{NL}$, and $\mathcal{L}^{Ref+NL}$, respectively. A Ref&NL data can be regarded as three input format data to yield three losses, while Ref-only and NL-only data can only compute the corresponding $\mathcal{L}^{Ref}$ and $\mathcal{L}^{NL}$. The final learning objective of UniCE is to minimize $\mathcal{L}^{Uni}$:
$\mathcal{L}^{Uni}=\mathcal{L}^{Ref}+\mathcal{L}^{NL}+\mathcal{L}^{Ref+NL},$ (9)
@@ 3 Evaluation
3.1 Experiment Setup
@@
Table 1: Statistics of datasets (part 1).
@@ 3.1.1 Datasets
Table 2: Statistics of datasets (part 2).
@@
APPS-Eval 0.3196 0.1814 0.1790 0.0315 0.0007 0.0011
HE-Eval - - 0.3022 - - 0.1499
To construct each code evaluation dataset, we first follow primitive NL and reference code in each corresponding base dataset. Then, for each paired NL and reference code in a code evaluation dataset, we generate an average of 20+ codes (generated from various LLMs, including CodeGen 350M&16B (Nijkamp et al., 2022), InCoder 1B&6B (Fried et al., 2022), and CodeX 13B&175B) (Chen et al., 2021) according to NL and additionally build an average of 100+ correct test cases according to reference code. To obtain these test cases, the following steps were implemented:
@@
Finally, we label each matched NL, reference code, and generated code by executing the generated code with all corresponding test cases to compute PassRatio via Eq. 1. Statistics of the datasets are presented in Table 1 and Table 2. As demonstrated in Table 1 and Table 2, there are notable disparities in the distributions of NL, RefCode (Reference Code), GenCode (Generated Code), and test cases across the three datasets.
3.1.2 Baselines

### v4

@@ Abstract.
Code Evaluation, Code Pre-trained Language Model, Code Generation.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Automatic evaluation of code generation is significant and promising in the fields of natural language processing (NLP) and software engineering. Due to the great potential of code generation in reducing development costs and revolutionizing programming modes, both industry and academia have devoted substantial attention to it (Li et al., 2022; Mukherjee et al., 2021; Yin and Neubig, 2018; Chen et al., 2021; Shen et al., 2022; Dong et al., 2023c). Code generation has achieved remarkable developments in the past few years (Fried et al., 2022; Nijkamp et al., 2023; Dong et al., 2023b; Jiang et al., 2023; Li et al., 2024b), but CEMs still need to catch up. It is challenging to evaluate the competitiveness of various approaches without proper CEM, which hampers the development of advanced techniques for code generation. A range of code generation subtasks would benefit from valid code evaluation, including code completion (Guo et al., 2022b; Lu et al., 2022), code translation (Rozière et al., 2020; Zhu et al., 2022), code search (Sun et al., 2022; Arakelyan et al., 2022), etc. Therefore, research on code evaluation is necessary and should be put on the agenda.
@@
Figure 1. Results of evaluating the generated code implementing bubble sort using different CEMs. BLEU and CodeBLEU score the truly functional correct code (c) lower than the incorrect code (b).
@@
Some commonly used match-based CEMs treat code as text, such as BLEU (Papineni et al., 2002) and Accuracy, which focus on basic and lexical-level features. They compute scores mainly based on n-gram co-occurrence statistics. CodeBLEU (Ren et al., 2020) additionally takes into account the structure of code, i.e., abstract syntax tree and data flow. However, the preceding CEMs have deficiencies in identifying code relationships, because code is mainly evaluated based on functional correctness rather than exact/fuzzy match to reference code, and match-based CEMs cannot account for the large and complex space of code functionally equivalent to reference code (Inala et al., 2022). For example, in Fig. 1, code (a) and code (b) have a much higher similarity of tokens or structures than code (c). However, through execution, we realize that code (a) and code (c) are different renderings of the same function. By contrast, the execution result of code (b) differs dramatically from both other codes, and code (b) even fails to compile. As a result, merely measuring the similarity of token/structure is insufficient for code evaluation.
@@
LLMs pre-trained on code have demonstrated outstanding results in code generation tasks (Chen et al., 2021; Fried et al., 2022; Li et al., 2022; Dong et al., 2023a, 2024a), which are fundamentally dependent on exceptional code comprehension. Excellent code comprehension is a crucial element for facilitating code evaluation. We hypothesize that LLMs pre-trained on code possess the ability to evaluate code. However, due to the training strategy of predicting the next token according to context, they lack awareness of evaluating code for functional correctness. Our objective is to instruct LLMs to evaluate code effectively in terms of functional correctness.
@@
Another issue that requires resolution is that the existing match-based CEMs are exclusively confined to the Ref-only (consider only reference code) input format. This restriction presents three inherent disadvantages. First, for any code generation task, the correct solutions are not finite, but rather, they are inexhaustible. In this context, the provided reference code merely represents one correct solution among a vast multitude. Therefore, it is overly narrow to compare the generated code solely with one correct solution. Second, they neglect the natural language (NL) description, which is a rich repository of information and a real requirement source. Third, these metrics are unusable in the absence of a reference code. This situation is quite commonplace in real-world evaluations where a correct solution is not always readily available. It is similar to code grading techniques in education, where grading often needs to be flexible and adaptable to different solutions that may not have a single correct answer. Therefore, expanding the input format of CEM is necessary.
@@
In this paper, we propose an effective LLM-based CEM, called CodeScore, which measures the functional correctness of generated codes on three input formats (Ref-only, NL-only, and Ref&NL). To obtain CodeScore, we present a code evaluation learning framework, UniCE, for tuning LLMs to estimate execution similarities with unified input. Specifically, we finetune LLMs to learn PassRatio and Executability of generated code, where Executability is devised to distinguish between compilation errors and output errors for code with PassRatio equal to 0. Generally, codes exhibiting higher functional correctness will pass more test cases, thereby achieving a higher PassRatio  [^1] Note that, although PassRatio varies across different test cases, it tends to yield a higher PassRatio for high-quality code, since we generate a large number of test cases. This phenomenon is somewhat akin to the process of human feedback. Despite the inherent variability in scores assigned by different human evaluators, the overarching trend remains consistent.. Consequently, for unexecutable codes, the model tends to assign scores approaching zero. In contrast, for codes demonstrating superior functional correctness, the model is likely to assign higher scores. CodeScore has the following advantages: 1) CodeScore has excellent evaluation performance, which achieves state-of-the-art performance correlation with functional correctness on multiple code evaluation datasets. 2) CodeScore provides three application scenarios (Ref-only, NL-only, and Ref&NL) for code evaluation with unified input, while traditional CEMs only consider Ref-only. Our major contributions can be summarized as follows:
@@
  • We propose an efficient and effective LLM-based CEM, CodeScore, that accommodates the functional correctness of generated codes from an execution viewpoint. [^2] https://huggingface.co/dz1/CodeScore
@@
  • CodeScore substantially outperforms match-based CEMs and LLM-based EMs, and achieves state-of-the-art performance on multiple code evaluation datasets.
2. Background & Related Work
In this section, we first introduce code generation, and then discuss code evaluation based on three types of EMs, including Match-based CEMs, Execution-based CEMs, and LLM-based EMs.
2.1. Code Generation
Code generation technology can automatically generate source code for software, achieving the purpose of machine-driven programming based on user requirements. Due to the rapid growth of code data and the continuous improvement of deep learning model capabilities, using deep learning for program generation has become the mainstream research direction (Raychev et al., 2014; Ling et al., 2016; Yin and Neubig, 2018; Wei et al., 2019; Sun et al., 2020; Mukherjee et al., 2021; Zhao et al., 2023; Jiang et al., 2024). In recent years, the rise of pre-training techniques has provided new momentum for code generation. For example, studies like CodeT5 (Wang et al., 2021) and UniXcoder (Guo et al., 2022a) pre-train models for completing code generation tasks. As the number of model parameters increases, researchers have observed the phenomenon of performance emergence in large language models (LLMs). . LLMs such as AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2023), WizardCoder (Luo et al., 2023), ChatGPT (OpenAI, 2023a), CodeGeeX (Zheng et al., 2023), Starcoder (Li et al., 2023), and CodeLlama (Rozière et al., 2023) have demonstrated promising code generation performance. Currently, code generation technology and tools have been widely adopted in software development, such as Copilot (Chen et al., 2021), significantly enhancing the efficiency of developers. Assessing the quality of generated code has remained a critical problem in the development of code generation technology, directly influencing its advancement and evolution.
2.2. Code Evaluation
Match-based CEMs.
Besides these commonly used BLEU (Papineni et al., 2002), Accuracy, and CodeBLEU (Ren et al., 2020), some niche CEMs (Popovic, 2015) are also applied to code evaluation, e.g., METEOR (Banerjee and Lavie, 2005), ROUGE (Lin, 2004), and CrystalBLEU (Eghbali and Pradel, 2022). However, these aforementioned match-based CEMs merely measure the surface-level differences in code and do not take into account the functional correctness of the generated code.
Execution-based CEMs.
They attempt to handle these issues by running tests for generated code to verify its functional correctness (Kulal et al., 2019; Hendrycks et al., 2021; Hao et al., 2022). However, they come with several caveats: 1) It assumes that test cases have been given and all dependencies have been resolved. For each code generation task, supplying adequate test cases is a burden in practice, and the dependencies required vary from task to task. 2) Enormous computational overhead needs to be afforded. All generated code requires execution separately for each corresponding test case, which leads to enormous CPU and I/O overhead. 3) Execution with isolation mechanisms. The generated code could have some security risks, such as deleting files on the disk or implanting computer viruses, especially if the training data of code generation models is attacked. In a word, they are usually costly, slow, and insecure, which are often unavailable or ineffective in real-world scenarios.
LLM-based EMs.
Effective evaluation of generated results is hard for both text and code generation. They likewise face the same issue of poor evaluation metrics (EMs). A recent popular trend in evaluating text generation is the design of automatic EMs based on LLMs. A part of LLM-based EMs (Rei et al., 2021; Wan et al., 2022; Rei et al., 2022) follows COMET (Rei et al., 2020) to learn high-quality human judgments of training data, which is a problem for code evaluation to obtain. Another part relies on LLM extracting token embeddings to calculate scores like BERTScore (Zhang et al., 2020), such as (Zhao et al., 2019; Sellam et al., 2020; Yuan et al., 2021; Reimers and Gurevych, 2019). A concurrent work named CodeBERTScore (Zhou et al., 2023) tries to use the same way as BERTScore with LLM pre-trained on code. However, they do not teach LLMs to learn code evaluation effectively, in other words, LLMs are still confused about how to evaluate code. Therefore, they exhibit suboptimal performance in code evaluation, as evidenced by our experimental results.
Figure 2. Examples of three input formats for code evaluation.
3. Methodology
@@ 3.1. CodeScore
(1) $\operatorname{PassRatio}=\frac{1}{|C_{p}|}\sum_{c\in C_{p}}\mathbb{I}\left\{\operatorname{Eval}\left(\mathbf{g}_{p},\mathcal{I}_{p,c}\right)=\mathcal{O}_{p,c}\right\}.$
where ||$\cdot$|| ⋅ | indicates the element number of a set, $\mathbb{I}\left\{\cdot\right\}$ is an indicator function, which outputs 1 if the condition is true and 0 otherwise, and $\operatorname{Eval}\left(\mathbf{g}_{p},\mathcal{I}_{p,c}\right)$ represents an evaluation function that obtains outputs of code $\mathbf{g}_{p}$ by way of executing it with $\mathcal{I}_{p,c}$ as input.
@@
(2) $\operatorname{Executability}=\left\{\begin{aligned} &1,\ if\ code\ is\ executable,\\ &0,\ otherwise.\end{aligned}\right.$
Figure 3. Diagram of UniCE, where the left side of the figure shows its model architecture, and the right side of the figure shows the example (case in Fig. 1) of input and output.
@@
(3) $(\operatorname{CodeScore},\operatorname{Exec})=\operatorname{UniCE}(\mathbf{x}),$
@@
We encourage UniCE to learn code execution (i.e., PassRatio and Executability) by minimizing loss function $\mathcal{L}$, which consists of two components:
(4) $\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{E},$
where $\mathcal{L}_{C}$ focuses on predicting PassRatio, and $\mathcal{L}_{E}$ on predicting code execution correctness. $\mathcal{L}_{C}$ and $\mathcal{L}_{E}$ are defined as:
(5) $\mathcal{L}_{C}=\left(\operatorname{CodeScore}-\operatorname{PassRatio}\right)^{2},$
(6) $\mathcal{L}_{E}=-\log\mathbf{p}(\operatorname{Exec}|\operatorname{Executability}),$
where $\mathcal{L}_{C}$ measures the squared difference between the predicted $\operatorname{CodeScore}$ and the actual PassRatio. $\mathcal{L}_{E}$ represents the negative log of the conditional probability of $\operatorname{Exec}$ given its Executability. The conditional probability is modeled as:
(7) $\mathbf{p}(\operatorname{Exec}|\operatorname{Executability})=\left\{\begin{aligned} &\mathbf{p}(\operatorname{Exec}),&\text{if }\operatorname{Executability}=1,\\ &1-\mathbf{p}(\operatorname{Exec}),&\text{otherwise},\end{aligned}\right.$
where $\mathbf{p}(\operatorname{Exec})$ is the predicted probability of successful execution.
3.2. UniCE
@@
3.2.1. Pooling Layer
For LLMs, the pooling layer plays a critical role in enhancing the model’s ability to capture and utilize information more effectively. The work (Tenney et al., 2019; Zhang et al., 2020; Rei et al., 2020) shows that exploiting information from different layers of LLM generally results in superior performance than only the last layer. Therefore, following the work (Peters et al., 2018), we pool information from different layers by using a layer-wise attention mechanism and the final embedding of a token $t$ can be computed as:
(8) $e_{t}=\gamma\sum^{l}_{k=1}e^{k}_{t}h^{k},$
@@ 3.2.1. Pooling Layer
3.2.2. Unified Embedding
We require an efficient and comprehensive representation to encapsulate the unified input sequence $x$. Generally, there are two standard methods to extract the representation of $x$, i.e., averaging all token embeddings and using the first token embedding. While the first method is straightforward and includes information from all tokens, it may dilute the significance of more critical tokens and introduce extraneous noise. The first token of our base models is specifically designed to be a summary token [^6] During the pre-training of our base models (such as CodeBert, GraphCodeBert, and UniXcoder), the first input token is typically the CLS token (short for ”classifier”), which enables the model to consider global contextual information during the encoding process through self-supervised learning methods. Therefore, the representation of this first token is usually used to represent the entire input sequence.. Moreover, the work (Ranasinghe et al., 2020; Wan et al., 2022) also proves the superiority of using the first token embedding compared to averaging all token embeddings in various applications. Thus, we employ the final embedding of the first token $e_{first}$ as the representation of the unified input sequence $x$.
3.2.3. Unified training
In UniCE, $e_{first}$ is fed to a feed-forward neural network to output a score and/or a category. To unify three evaluation input formats into UniCE, we apply multi-task learning for training. Specifically, for each step, we assign three sub-steps for three input formats, yielding $\mathcal{L}^{Ref}$, $\mathcal{L}^{NL}$, and $\mathcal{L}^{Ref+NL}$, respectively. A Ref&NL data can be regarded as three input format data to yield three losses, while Ref-only and NL-only data can only compute the corresponding $\mathcal{L}^{Ref}$ and $\mathcal{L}^{NL}$. The final learning objective of UniCE is to minimize $\mathcal{L}^{Uni}$:
(9) $\mathcal{L}^{Uni}=\mathcal{L}^{Ref}+\mathcal{L}^{NL}+\mathcal{L}^{Ref+NL},$
@@ 4. Evaluation
Our five RQs aim to evaluate the efficacy and practicality of our approach compared to existing EMs. RQ1 and RQ4 assess our approach against current EMs through experiments and human evaluations, ensuring a comprehensive analysis from both quantitative and qualitative perspectives. RQ2 and RQ3 involve ablation studies to pinpoint the individual and combined impacts of our approach’s main components. RQ5 evaluates our approach’s real-world applicability through case studies.
4.1. Experiment Setup
@@
Table 1. Statistics of datasets (part 1).
@@ 4.1.1. Datasets
Table 2. Statistics of datasets (part 2).
@@
HE-Eval - - 0.3695 - - 0.1591
To construct each code evaluation dataset, we first follow primitive NL and reference code in each corresponding base dataset. Then, for each paired NL and reference code in a code evaluation dataset, we generate an average of 20+ codes (generated from various LLMs, including CodeGen 350M&16B (Nijkamp et al., 2023), InCoder 1B&6B (Fried et al., 2022), and CodeX 13B&175B (Chen et al., 2021). For HE-Eval dataset, we also consider the latest state-of-the-art LLMs including StarCoder 15.5B (Li et al., 2023), CodeLlama 34B (Rozière et al., 2023), and GPT-4 (OpenAI, 2023b) besides the aforementioned LLMs.) according to NL and additionally build an average of 100+ correct test cases according to reference code. To obtain these test cases, the following steps were implemented:
@@
Finally, we label each matched NL, reference code, and generated code by executing the generated code with all corresponding test cases to compute PassRatio via Eq. 1. Statistics of the datasets are presented in Table 1 and Table 2777For each generated code, we employ extended test cases of the corresponding task to compute its PassRatio and Passability. We compute the average number of PassRatio and Passability, i.e., AvgPassRatio and Pass@1, on the train, dev, and test sets of each dataset and display them in Table 2.. As demonstrated in Table 1 and Table 2, there are notable disparities in the distributions of NL, RefCode (Reference Code), GenCode (Generated Code), and test cases across the three datasets. Specifically,
  • APPS-Eval has 267,162 training examples and 33,395 examples each for dev and test sets. Each task typically includes 1 NL, 13 RefCode, and 42 GenCode, with average token lengths of 263.8 for NL, 86.3 for RefCode, and 76.8 for GenCode. Extended test cases average 181 per task, compared to the original 13. The AvgPassRatio for train, dev, and test sets are 0.3196, 0.1814, and 0.1790, respectively, while Pass@1 are 0.0315, 0.0007, and 0.0011, respectively.
  • MBPP-Eval has 15,679 training examples and 3,000 examples each for dev and test sets. Each task typically includes 1 NL, 1 RefCode, and 24 GenCode, with average token lengths of 15.5 for NL, 32.5 for RefCode, and 26.7 for GenCode. Extended test cases average 102 per task, compared to the original 3. The AvgPassRatio for train, dev, and test sets are 0.2832, 0.2571, and 0.2890, respectively, while Pass@1 are 0.0674, 0.0494, and 0.0760, respectively.
  • HE-Eval has 4,221 test examples. Each task typically includes 1 NL, 1 RefCode, and 26 GenCode, with average token lengths of 61.9 for NL, 24.4 for RefCode, and 41.6 for GenCode. Extended test cases average 108 per task, compared to the original 8. The AvgPassRatio for the test set is 0.3695, while Pass@1 is 0.1591.
4.1.2. Baselines